Connect Quiz (Chapter 4) Post-Class Calculate the following values. 3 903A) tos a. 10! 6! b. 10C6 c. 10P6 00:24:32
Answer #1

a)10! =10*9*8*7*6*5*4*3*2*1=3628800

b) 6! =720

c)10C6 =10!/(6!*4!)=3628800/(720*24)=210
d)10P6 =10!/4! =3628800/24=151200